What is the equation in slope intercept form with the points (5,1) and (10,10)
Answer by jim_thompson5910(35256)


Slope:
m = (y2 - y1)/(x2 - x1)
m = (10 - 1)/(10 - 5)
m = 9/5

Point Slope form:
y - y1 = m(x - x1)
y - y1 = (9/5)(x - x1) ..... plug in the slope found earlier
y - 1 = (9/5)(x - 5) .... plug in the point (5,1)
y - 1 = (9/5)x + (9/5)(-5)
y - 1 = (9/5)x - 9
y - 1+1 = (9/5)x - 9+1 .... add 1 to both sides
y = (9/5)x - 8

Answer: y = (9/5)x - 8